Archive for the ‘Incidents’ Category

Incident: columbo.core

Posted: Thursday, December 1st, 2016 at 09:35 by Richard Partridge

We are currently experiencing a loss of connectivity to columbo.core. Engineers are already investigating and we are organising local engineering resources to attend site.

Further updates will be provided as and when they become available.

  1. Richard Partridge says:

    We have regained connectivity to the core router and initial evidence suggests this loss of connectivity relates to a power failure to our cabinet. We are continuing our investigation and the site should be considered at risk for the time being.

  2. Ryan Berrisford says:

    Although power has been restored to the Columbo core router we are still seeing some services as down. Engineers are continuing to investigate and further updates will follow.

  3. Ryan Berrisford says:

    Services have now been fully restored to both DSL and Ethernet services. We are continuing to monitor and a full incident report will follow in due course.

  4. Ryan Berrisford says:

    An incident report for this issue is now available here.

Incident: Nottingham WBC

Posted: Thursday, November 10th, 2016 at 15:21 by Richard Partridge

Engineers are currently investigating an issue causing a loss of service for DSL connections served by the Nottingham WBC node.

Further updates will be provided as and when they become available.

 

  1. Richard Partridge says:

    Our engineers have corrected the problem and service is now restored.

    We apologise for any inconvenience caused.

  2. Neil Watson says:

    A report is now available for this incident which can be found here. Apologies for the problems that this loss of service will have caused for our customers.

Incident: Telehouse East Satellite

Posted: Tuesday, November 8th, 2016 at 11:16 by Richard Partridge

We are currently aware of an issue affecting one of our satellite devices in Telehouse East causing a loss of service for a small number of customers.

Engineers are already investigating and a further engineer is on his way to provide on-site assistance.

  1. Richard Partridge says:

    Our engineer is now on site and investigating.

    Further information will be provided as and when it becomes available.

  2. Richard Partridge says:

    A hardware fault has been rectified and service is now restored.

    We apologise for any inconvenience caused.

  3. Jake Turner says:

    While most services connected to this edge device have had service restored, we are still seeing residual issues with a very small number of customers.

    An upgrade and reboot of the device will be carried out at 23:00 this evening. The emergency maintenance should take no longer than 1 hour however actual downtime for each customer is likely to be around 5 minutes.

  4. Jake Turner says:

    This work is now complete and service has been restored to the affected customers.

  5. Ryan Berrisford says:

    An incident report for this issue is now available here.

Incident: managed-lns01.inx.dsl

Posted: Wednesday, November 2nd, 2016 at 18:25 by Mark Yardley

We have recently experienced an unscheduled reload of managed-lns01.inx.dsl serving our managed EWCS platform. This will have caused a brief interruption of service to customers connected to this particular LNS. Engineers are monitoring the device and will be investigating further to ascertain the cause.

We apologise for the inconvenience this may have caused.

Incident: VoIP Disruption

Posted: Wednesday, November 2nd, 2016 at 15:45 by Richard Partridge

One of our upstream carriers is experiencing difficulties which is having an adverse affect on outbound call setups. Engineers are working to bypass this problem and direct outbound calls towards other carriers.

Further updates will be provided as and when they become available.

  1. Richard Partridge says:

    The affected carrier has now been bypassed.

  2. Richard Partridge says:

    The upstream carrier has acknowledged the problem and is working on a resolution. We will continue to re-route calls around this problem until a solution has been put in place.

  3. David Derrick says:

    The affected carrier resolved their problem and has now been brought back into service.

Incident: Broadband Connections

Posted: Saturday, October 22nd, 2016 at 13:36 by Neil Watson

We are aware of an incident earlier today that affected a proportion of our broadband base which prevented them from being able to connect to our service for around 15 – 30 minutes. All services have been restored and we continue to monitor to ensure that all customers can connect as expected. We will investigate further and will post a report in due course. Apologies to any customer that was impacted.

  1. Adam Heath says:

    An incident report for this issue is now available here

Incident: Vodafone Circuits

Posted: Monday, October 17th, 2016 at 16:47 by Adam Heath

We are currently aware of an issue affecting a number of Vodafone based circuits.  We are working with our supplier who are currently investigating.

We will publish further updates as they become available.

  1. Adam Heath says:

    After initial testing Vodafone have acknowledged the issue on their side. Further updates will be available shortly.

  2. Ryan Berrisford says:

    Vodafone have confirmed that following diagnostics, engineers have been dispatched to Sovereign House as an issue is suspected within that domain. Estimated time of arrival is currently 21:00.

    Further updates will follow once the issue has been further isolated or resolved.

  3. Ryan Berrisford says:

    Vodafone have confirmed that although engineering resource has reached site, there was a delay in arrival due to the black cab protest taking place in Central London. Fault resolution has therefore been delayed slightly but further updates will be made available shortly.

  4. Ryan Berrisford says:

    Vodafone have advised that the fault is still in hand with engineering resource at Telecity, Soveriegn House. We are pressing for resolution and updates and will post further information as it becomes available.

  5. Ryan Berrisford says:

    Service has now been restored to the affected circuits. We are awaiting full resolution details and will update this post in due course.

    We apologise for the inconvenience caused by this incident.

  6. Ryan Berrisford says:

    Vodafone have confirmed the incident was caused by an issue with the line card at Telecity, Sovereign House. The issue is confirmed as resolved by Vodafone.

    We apologise for the inconvenience caused during this incident.

Incident: Faraday

Posted: Wednesday, October 12th, 2016 at 12:56 by Jake Turner

We are aware of an issue affecting customers on our Faraday core. Engineers are investigating and further updates will be provided when available.

  1. Jake Turner says:

    Service has been restored. Engineers will continue to monitor and provide further information when available.

  2. Richard Partridge says:

    It appears that there are still residual problems affecting service for a small number of customers. These are being investigated on a case by case basis.

    We apologise for the inconvenience caused and are working as quickly as possible to resolve the issue. Further updates will be provided as and when they become available.

  3. Richard Partridge says:

    Engineers are continuing to work on the problem and the affected circuits are being brought back into service one at a time.

    Further updates will be provided shortly.

  4. Phil Davies says:

    More circuits have now been successfully restored and engineers are continuing to work on the few remaining which are down.

    Further updates to follow.

  5. Richard Partridge says:

    With only a few circuits remaining with issues, customers affected will now be contacted directly with any further updates.

    Thank you for your patience.

Incident: Sheffield

Posted: Friday, September 9th, 2016 at 17:12 by Mark Yardley

We are aware of an issue affecting customers on our Sheffield core. Engineers are investigating and further updates will be provided when available.

  1. Jake Turner says:

    Service has been restored. Engineers will continue to monitor and provide further information when available.

Ongoing VoIP traffic issues

Posted: Tuesday, August 16th, 2016 at 17:52 by Neil Watson

The monitoring as advised previously below has identified specific actions for us to take to overcome the issues impacting a relatively small number of customers using VoIP on a subset of the six nodes previously identified. We will implement these actions during the course of the evening and the pre-published maintenance window tonight. Users may see their service reconnect during the window. We anticipate that this will be an instant reconnect, however we would ask customers to reboot their equipment should subsequent connection problems be experienced. We will provide an update once normal daytime-business traffic levels are reached.

  1. Neil Watson says:

    As a result of the work carried out last night, we are seeing greatly improved performance on a number nodes. Work continues to correct the problem for the small amount of other customers that remain affected. As well as aiming to correct the root cause, we are also attempting to put a work-around in place to alleviate the symptoms. We anticipate that this will be in place today, however will confirm when complete.

  2. Neil Watson says:

    We have continued to work on the traffic management platform throughout the day with the assistance of the manufacturer to resolve the issues that continue to be seen. This work has so far not been successful; however we have implemented a workaround which has significantly improved the performance of VoIP on one test node.

    We have also been monitoring the changes made last night to a number of nodes and have seen no further issues reported today. We will therefore replicate those actions tonight during the pre-published maintenance window on other nodes. Users may see their service reconnect during the window. We anticipate that this will be an instant reconnect, however we would ask customers to reboot their equipment should subsequent connection problems be experienced. We will continue to investigate the root cause of the issue and will make further updates.

  3. Neil Watson says:

    Monitoring this morning has proved positive and we believe that last night’s actions have restored service quality to customers using VoIP. We are continuing to investigate the root cause and will update with more information as soon as we are able. If any customer experiences further issues please contact our support team.

  4. Neil Watson says:

    Following further final work last night and a period of monitoring this morning, we are now able to confirm that the VoIP issues have been resolved for all customers. We apologise for the extended problems that have been experienced.